The sum of a number is it's reciprocal is 65/8 determine the number
Answer by ewatrrr(24785)

n + 1/n = 65/8 |Multiplying thru by 8n so as all denominators = 1
8n^2 -65n + 8 = 0
(8n - 1)(n - 8) = 0, n = 1/8 0r 8